OpenStudy (anonymous):

Ok here's a Calc. Question. I understand how to do this but keep getting the wrong answer for my homework. Please help.... f(x)= (x-8) all over ((x-1)(x+1)) Use interval notation to indicate where f(x) is continuous.

OpenStudy (science0229):

First, what are the x-values that make the denominator 0?

OpenStudy (anonymous):

1 and -1.

OpenStudy (science0229):

Right, so if we were to graph this, those would be the 2 asymptotes.

OpenStudy (anonymous):

yeah that's the graph that I got was asymptotes

OpenStudy (science0229):

Now, is there any y-asymptotes?

OpenStudy (anonymous):

no.....

OpenStudy (science0229):

are you sure?

OpenStudy (science0229):

\[\lim_{x \rightarrow \pm \infty}\frac{ x-8 }{ x^2-1 }=\lim_{x \rightarrow \pm \infty}\frac{ \frac{ 1 }{ x }-\frac{ 8 }{ x^2 } }{ 1-\frac{ 1 }{ x^2 } }=0\]

OpenStudy (science0229):

As x approaches positive/negative infinity, the function approaches 0.

OpenStudy (science0229):

Now, is there any x-intercepts?

OpenStudy (science0229):

@jclark1989

OpenStudy (science0229):

BTW we're doing this to graph the function because once you graph it, it is very easy to see where the function is continuous.

OpenStudy (anonymous):

yeah.. the x intercepts would be 1 and -1 right?

OpenStudy (science0229):

Remeber that x intercepts are the points where the y-coordinate is 0.

OpenStudy (science0229):

In other words, f(x)=0

OpenStudy (anonymous):

ahh ok.

OpenStudy (science0229):

What did you get?

OpenStudy (science0229):

\[\frac{ x-8 }{ x^2-1 }=0\rightarrow x=8\]

OpenStudy (anonymous):

ahh. so the x cord. is 1 and -1.

OpenStudy (anonymous):

this graph is throwing me off. the y cord is 8?